what is the degree of unsaturation is a compound with molecular formula C9H2N4

Grade:11

2 Answers

Anish
2 years ago
Calculate the unsaturation number for your compound, which is now in the form CnHm, using this formula Ω = n - (m/2) + 1, where "Ω" is the degree of unsaturation--the number of pi bonds your compound has. For the example compound, C6H4, do this as follows: Ω = 6 - (4/2) + 1 = 6 - 2 + 1 = 5.
Unicorn
2 years ago
Remember this amazing formula which contains all elements that a compound conatain to calculate degree of unsaturation 
 
C(a) H(b) D(c) T(d) X(e) O(f) S(g) N(h) P(i)
Capital letters represents elements carbon hydrogen ....etc and in bracket represents no. Of atoms like C9H2N4 ((a=9,b=2,h=4))
 
Degree of unsaturation= (2a+2-[b+c+d+e] +h+i ) ÷2
 
Important point there is no contribution of oxygen and sulphur 
 
Hope it helps 
Approve:)